OpenStudy (anonymous):

What is the number of real solutions? -11x^2 = x+11 a.cannot be determined b.one solution c.two solutions d.no real solutions

OpenStudy (ranga):

Put the equation in standard form by adding -11x^2 to both sides. Then find the discriminant. If the quadratic equation is ax^2 + bx + c = 0 the discriminant D = b^2 - 4ac Find D and we can take the next step.

OpenStudy (ranga):

-11x^2 = x+11 Put the equation in standard form by adding 11x^2 to both sides. 0 = 11x^2 + x + 11 or 11x^2 + x + 11 = 0 Compare this to the standard quadratic equation ax^2 + bx + c = 0 What is a, b, c? What is discriminant = b^2 - 4ac ?